OpenStudy (anonymous):

The area of a square garden is 200 m². How long is the diagonal?

OpenStudy (anonymous):

wow that was wrong

OpenStudy (anonymous):

side is \[\sqrt{200}\] so diagonal is \[\sqrt{200}\times \sqrt{2}=\sqrt{400}=20\]

OpenStudy (kainui):

If you know the area of a square, you know that the square root of the area is the length of one side. The diagonal is just the hypotenuse of a right triangle, right? So you can now just apply the pythagorean theorem!

OpenStudy (anonymous):

Diagonal of square = sqrt2 * side

OpenStudy (anonymous):

20